# March 19, 2024

### v6

@@
Projective quantum measurement is a theoretically accepted process in modern quantum mechanics. However, its projection hypothesis is widely regarded as an experimentally established empirical law. In this article, we combine a previous result regarding the realization of a Hamiltonian process of the projection hypothesis in projective quantum measurement, where the complete set of the orbital observables of the center of mass of a macroscopic quantum mechanical system is restricted to a set of mutually commuting classical observables, and a previous result regarding the work required for an event reading (i.e., the informatical process in projective quantum measurement). Then, a quantum thermodynamic scheme is proposed for experimentally testing these two mutually independent theoretical results of projective quantum measurement simultaneously.
Keywords: quantum measurement theory, projection hypothesis, quantum thermodynamics, quantum Jarzynski equality
@@ 1 Introduction
In this article, we combine two mutually independent theoretical results of projective quantum measurement. The first, from Ref. [6], regards the realization of a Hamiltonian process of the projection hypothesis in projective quantum measurement, where the complete set of the orbital observables of the center of mass of a macroscopic quantum mechanical system is restricted to a set of mutually commuting classical observables [7, 8, 9]. The second, from Ref. [10], regards the work required for an event reading (i.e., the informatical process in projective quantum measurement). We also propose a quantum thermodynamic scheme to experimentally test these two theoretical results simultaneously.
Throughout this article, to model projective quantum measurement, we adopt the ensemble interpretation of quantum mechanics [9]. Specifically, for a given quantum system, a pure ensemble means an ensemble of copies of a state vector, and a non-trivial mixed ensemble means a statistical mixture of copies of at least two distinct state vectors. Supposing an eigenbasis of the state space (we take it as the eigenbasis of the discrete measured observable), we refer to an ensemble as a coherent ensemble when its elements have quantum coherence (i.e., are quantum superpositions of at least two distinct eigenstates) and as a classical ensemble when its elements do not have any quantum coherence (i.e., are eigenstates).
Projective quantum measurement has two temporally successive processes: non-selective measurement and its subsequent event reading. [^2] Intuitively, the term non-selective measurement means that, after only this dynamical process, events are in a mixture and no particular event is yet singled out. Non-selective measurement (also called a complete decoherence process or a quantum-to-classical transition [11]) means a dynamical process from a coherent pure state (that is, the state to be measured) to its classical mixed state (i.e., its diagonal part as a mixture of eigenstates). Event reading means an informatical process from a classical mixed state to a classical pure state (i.e., an eigenstate). We call the substance of the latter process the projection hypothesis (for its older form, see Refs. [7] and [12]).
@@
Here, we identify the novelty of the present work within the existing literature on quantum thermodynamics. At present, the energetics of projective quantum measurement is based on the quantum thermodynamics of information [1, 3, 19]. In this theory, the fundamental energy cost is considered for projective quantum measurement, which is done in the general positive operator-valued measure (POVM) measurement [4], in the ensemble description [1, 2] and the reset process of the measurement outcome stored in the memory system [19]. On the other hand, the present article and Refs. [10] and [13] consider the work required for event reading, that is, the informatical process in projective quantum measurement, which is independent from the former energy cost.
Throughout this article, we denote operators with a hat.
The rest of this article is organized as follows. In Sec. 2, we explain the systems to be considered, the treatment of a macroscopic quantum system, and the von Neumann-type interaction. Sec. 3 is devoted to a new and clear elaboration on the origin of the required work for event reading derived in Ref. [10]. In Sec. 4, we model the quantum thermodynamic scheme, to experimentally test the previously obtained mutually independent theoretical results of projective quantum measurement simultaneously, and clarify the classical nature of the required work for event reading. In Sec. 5, we conclude the article. In Appendices A and B, we present supplementary calculations. In Appendix C, we clarify the experimental platform for realizing the systems which appear in the quantum thermodynamic scheme.
@@ 2 Preliminaries
In this article, using the notation of Ref. [6] [^3] However, we denote $\Psi$ in Ref. [6] by $M$ here., we consider four systems in two groups:
@@
for a positive-valued c-number constant $\Lambda$ by the passive means of spatial translation of $A$ in quantum mechanics. [^4] For accounts for the passive means of spatial translation in quantum mechanics, see Chapter 6 of Ref. [22]. Usually, we assume that $\Lambda$ is strong enough to ignore the free Hamiltonians of $S$ and $A$ relatively to this interaction.
@@ 3 Work required for event reading
Here, we assume that the combined measured system $S$ is in a non-equilibrium and controlled redefined Hamiltonian process and is initially in thermal equilibrium at temperature $T$ [23]. For simplicity, we further assume that the combined measured system $S$ is weakly coupled to a canonical heat bath initially at temperature $T$ [^5] For the treatment of strong coupling between the combined measured system $S$ and a canonical heat bath, see Ref. [24].
@@
[…]
@@
For the entropy production $\sigma:=-\ln\rho$ attributable to the decay of the population of $\rho$ in the enlarged ensemble, Eq. (9) is analogous to a relaxation process in the conventional ensemble at $t=dt$
@@
In the statistical description, the self-similar relaxation means the entropy production $\sigma$ of 1 nat with respect to the systems without relaxation in the enlarged ensemble. This entropy production is attributable to the fact that, in the statistical description, the relaxation is in a cut-off stochastic process, and thus its occurrence is not mechanical but informative as opposed to the direct description.
@@
As shown in the next section, this information with respect to the systems without relaxation in the enlarged ensemble is lost by the combined measuring system $M$ and is acquired by the combined measured system $S$ just before event reading by $M$.
@@ 4 Quantum thermodynamic scheme
    which is a classical mixed state. [^8] In Ref. [10], we adopt the continuous superselection rule of $P^{a}$. So, there, the original Hamiltonian is compatible with this superselection rule.
@@
in the trigger of the event reading by $\psi$. Here, $\delta_{\mu}\Xi_{A}=\mu\dot{\Xi}_{A}$ is a spatial displacement of the origin of the rearranged coordinate system $\Xi_{A}^{a}$ of $A$ in the von Neumann-type interaction [^9] Here, we adopt the interaction picture of $M$, and we rescale time so that $\Lambda$ in Ref. [6] is unity. between $\psi$ and $A$ occurring over an elapsed time $\mu$ dominantly [6]. We can remove the dynamical degrees of freedom of $M$ from the whole system without changing the state of $S$ at the end of step (IV) if $S$ is isolated from $M$ after the inverse unitary transformation of step (IV) of the whole system (as a canonical transformation in quantum mechanics) [10, 13]; note that this unitary transformation does not change the state of $S$. In this case, time evolution of the state of $S$ in step (V) is given by (for its derivation, see Appendix B)
@@
for the variable $W$ of an amount of single-shot work [^10] Single-shot work is deterministic and has no fluctuations. on $S$ [23, 28] and the controlled redefined Hamiltonian ^$\widehat{{\cal H}}^{S}_{H,\lambda_{t}}$ of $S$ in the Heisenberg picture with a control parameter $\lambda_{t}$, which is further parameterized by time $t$ in $[0,t_{f}]$ according to the quantum force protocol [23]. Here, ${\cal T}_{>}$ is the time ordering of products and we set $\beta=1/(k_{B}T)$, $Z_{\lambda}={\rm tr}_{S}\bigl{[}e^{-\beta\widehat{{\cal H}}_{H,\lambda}^{S}}\bigr{]}$, and $t_{n}=nt_{f}/N$. This equality is
@@
    The operator to be averaged in the quantum Jarzynski equality (21) is defined over $t_{f}\geq t\geq t^{(0)}$ at an arbitrary intermediate time $t^{(0)}$ in the whole process $[0,t_{f}]$. [^11] For its derivation, see Appendix C.1 of Ref. [10]. Thus, the operator to be averaged is redefined by multiplying it by the factor $e^{-3\sigma}$.
@@
In quantum thermodynamics, the quantum work done by a work reservoir to a thermodynamic quantum mechanical system [^12] In modern thermodynamics, the thermodynamic systems and the systems in the thermodynamic limit are distinct classes of systems: a microscopic system in contact with a large canonical heat bath is a thermodynamic system. and the heat absorption from a heat bath by the system are distinguished as the change of the energy eigenvalues (the quantum Hamiltonian) of the system and the change of the corresponding occupation probabilities of the system, respectively, in the internal energy of the system [27, 30, 31].
@@ 5 Conclusion
Finally, to conclude this article, we explain how this result can be positioned within the existing literature on quantum thermodynamics from the perspective of the information-theoretical origins of work, in the setting of the quantum Jarzynski equality, in terms of the density matrix. The explanation consists of three stages.
@@
for the average quantum work $\overline{W}$ done by a work reservoir to a thermodynamic quantum mechanical system, which is in contact with a large canonical heat bath at temperature $T$, and the change in the non-equilibrium free energy $\Delta F$ of the system [^13] Non-equilibrium free energy is obtained from the equilibrium free energy by replacing the thermodynamic entropy with information-theoretical entropy [32]. induced by the work in $\overline{W}$. In general, the work $W$ fluctuates thermally and quantum mechanically. In the thermodynamic limit of the system, the thermal fluctuations in the system become negligible. The change in the non-equilibrium free energy $\Delta F$ of the system has an information-theoretical origin in terms of the density matrix:
@@
where $\widehat{\varrho}$ is the terminal quantum state of the system and $\widehat{\varrho}_{\rm can}$ is the initially prepared canonical thermal equilibrium quantum state of the system at temperature $T$ [33]. Here, the relative von Neumann entropy is always non-negative: $\Delta F\geq 0$. It should be noted that this expression (29) and (30) for the work was extended, in the single-shot work, to the Gibbs-preserving and completely-positive and trace-preserving map of the system in the context of quantum resource theory, where the extractable work and the work for state formation are each individually regarded as a resource and the thermodynamic limit of the system is unnecessary [32, 34, 35, 36, 37].
@@
Here, $W_{\rm e.r.}$ is internal work. Namely, in the composite system of systems $S$ and $M$, the work for $S$ and that for $M$ cancel each other out and there is no net extractable work and no net energy cost.
We now conclude this article. As seen from the above explanations of the information-theoretical origins of work in quantum thermodynamics in three stages, the extractable quantum work $-W_{\rm f.b.}$ characterizes the information-to-energy conversion (see Ref. [41] for a renowned experiment in the classical mechanical context), but the classical single-shot work $W_{\rm e.r.}$ characterizes the physicality of the projection hypothesis in the setting of the quantum Jarzynski equality. Namely, in the ensemble interpretation of quantum mechanics, we have theoretically shown a way to experimentally examine the physicality of the projection hypothesis [13]. This is the implication of the main results in the fields of quantum information and quantum thermodynamics.
@@ Appendix C Experimental platform
In this appendix, we clarify the experimental platform for realizing the four systems which appear in the quantum thermodynamic scheme proposed in Sec. 4. A model of the quantum measurement steps in the proposed scheme is shown in Fig. C1.
Figure C1: Quantum measurement steps in the proposed quantum thermodynamic scheme. There are three steps: first, the non-selective measurement (n.s.m.) in the composite system $S=S_{0}+A^{\prime}$; second, the entangling (ent.) process between systems $S_{0}$ in $S$ and $\psi$ in $M$; and third, the von Neumann-type interaction (v.N. int.) between systems $\psi$ and $A$. Each dashed arrow from system $X$ to system $Y$ indicates the interaction process (e.g., the von Neumann-type interaction) between systems $X$ and $Y$, where the state of system $X$ does not change.
In Fig. C1, $S_{0}$ and $\psi$ are quantum mechanical measured and event-reading systems, respectively. In the context of quantum measurement theory [9], they are distinguished from each other. See also remark (E).

### v8

# June 20, 2024

@@
Projective quantum measurement is a theoretically accepted process in modern quantum mechanics. However, its projection hypothesis is widely regarded as an experimentally established empirical law. In this paper, we combine a previous result regarding the realization of a Hamiltonian process of the projection hypothesis in projective quantum measurement, where the complete set of the orbital observables of the center of mass of a macroscopic quantum mechanical system is restricted to a set of mutually commuting classical observables, and a previous result regarding the work required for an event reading (i.e., the informatical process in projective quantum measurement). Then, a quantum thermodynamic scheme is proposed for experimentally testing these two mutually independent theoretical results of projective quantum measurement simultaneously.
Keywords: Quantum measurement theory; projection hypothesis; quantum thermodynamics; quantum Jarzynski equality.
@@ 1 Introduction
In this paper, we combine two mutually independent theoretical results of projective quantum measurement. The first, from Ref. [6], regards the realization of a Hamiltonian process of the projection hypothesis in projective quantum measurement, where the complete set of the orbital observables of the center of mass of a macroscopic quantum mechanical system is restricted to a set of mutually commuting classical observables [7, 8, 9]. The second, from Ref. [10], regards the work required for an event reading (i.e., the informatical process in projective quantum measurement). We also propose a quantum thermodynamic scheme to experimentally test these two theoretical results simultaneously.
Throughout this paper, to model projective quantum measurement, we adopt the ensemble interpretation of quantum mechanics [9]. Specifically, for a given quantum system, a pure ensemble means an ensemble of copies of a state vector, and a non-trivial mixed ensemble means a statistical mixture of copies of at least two distinct state vectors. Supposing an eigenbasis of the state space (we take it as the eigenbasis of the discrete measured observable), we refer to an ensemble as a coherent ensemble when its elements have quantum coherence (i.e., are quantum superpositions of at least two distinct eigenstates) and as a classical ensemble when its elements do not have any quantum coherence (i.e., are eigenstates).
Projective quantum measurement has two temporally successive processes: non-selective measurement and its subsequent event reading.aaaIntuitively, the term non-selective measurement means that, after only this dynamical process, events are in a mixture and no particular event is yet singled out. Non-selective measurement (also called a complete decoherence process or a quantum-to-classical transition [11]) means a dynamical process from a coherent pure state (that is, the state to be measured) to its classical mixed state (i.e., its diagonal part as a mixture of eigenstates). Event reading means an informatical process from a classical mixed state to a classical pure state (i.e., an eigenstate). We call the substance of the latter process the projection hypothesis (for its older form, see Refs. [7] and [12]).
@@
Here, we identify the novelty of the present work within the existing literature on quantum thermodynamics. At present, the energetics of projective quantum measurement is based on the quantum thermodynamics of information [1, 3, 19]. In this theory, the fundamental energy cost is considered for projective quantum measurement, which is done in the general positive operator-valued measure (POVM) measurement [4], in the ensemble description [1, 2] and the reset process of the measurement outcome stored in the memory system [19]. On the other hand, the present paper and Refs. [10] and [13] consider the work required for event reading, that is, the informatical process in projective quantum measurement, which is independent from the former energy cost.
Throughout this paper, we denote operators with a hat.
The rest of this paper is organized as follows. In Sec. 2, we explain the systems to be considered, the treatment of a macroscopic quantum system, and the von Neumann-type interaction. Section 3 is devoted to a new and clear elaboration on the origin of the required work for event reading derived in Ref. [10]. In Sec. 4, we model the quantum thermodynamic scheme, to experimentally test the previously obtained mutually independent theoretical results of projective quantum measurement simultaneously, and clarify the classical nature of the required work for event reading. In Sec. 5, we conclude the paper. In Appendices A and B, we present supplementary calculations. In Appendix C, we clarify the experimental platform for realizing the systems which appear in the quantum thermodynamic scheme.
@@ 2 Preliminaries
In this paper, using the notation of Ref. [6]bbbHowever, we denote $\Psi$ in Ref. [6] by $M$ here., we consider four systems in two groups:
@@
for a positive-valued c-number constant $\Lambda$ by the passive means of spatial translation of $A$ in quantum mechanics.cccFor accounts for the passive means of spatial translation in quantum mechanics, see Chapter 6 of Ref. [22]. Usually, we assume that $\Lambda$ is strong enough to ignore the free Hamiltonians of $S$ and $A$ relatively to this interaction.
@@ 3 Work Required for Event Reading
Here, we assume that the combined measured system $S$ is in a non-equilibrium and controlled redefined Hamiltonian process and is initially in thermal equilibrium at temperature $T$ and until then in thermal contact with a heat bath at temperature $T$ [23].
@@
Let $\rho=\rho(t)$ denote the statistical weight of the system without relaxation in the ensemble of copies of the system at a given time $t$. We set $\rho(0)=1$.dddWhen we identify $\rho$ with the normalization factor of a density matrix, the relaxation process is truncated from the time evolution of the density matrix. See Appendix A.1.
@@
For the entropy production $\sigma:=-\ln\rho$ attributable to the decay of $\rho$ from unity, Eq. (9) is analogous to a relaxation process in the conventional ensemble at $t=dt$
@@
In the statistical description, the self-similar relaxation means the entropy production $\sigma$ of 1 nat with respect to the statistical weight of the system without relaxation in the enlarged ensemble. This entropy production is attributable to the fact that, in the statistical description, the relaxation is in a cut-off stochastic process, and thus its occurrence is not mechanical but informative as opposed to the direct description.
@@
As shown in the next section, this 1 nat of information with respect to the statistical weight of the system without relaxation in the enlarged ensemble is lost by the combined measuring system $M$ and is acquired by the combined measured system $S$ just before event reading by $M$.
@@ 4 Quantum Thermodynamic Scheme
    which is a classical mixed state.fffIn Ref. [10], we adopt the continuous superselection rule of $P^{a}$. So, there, the original Hamiltonian is compatible with this superselection rule.
@@
in the trigger of the event reading by $\psi$. Here, $\delta_{\mu}\Xi_{A}=\mu\dot{\Xi}_{A}$ is a spatial displacement of the origin of the rearranged coordinate system $\Xi_{A}^{a}$ of $A$ in the von Neumann-type interactiongggHere, we adopt the interaction picture of $M$, and we rescale time so that $\Lambda$ in Ref. [6] is unity. between $\psi$ and $A$ occurring over an elapsed time $\mu$ dominantly [6]. We can remove the dynamical degrees of freedom of $M$ from the whole system without changing the state of $S$ at the end of step (IV) if $S$ is isolated from $M$ after the inverse unitary transformation of step (IV) of the whole system (as a canonical transformation in quantum mechanics) [10, 13]; note that this unitary transformation does not change the state of $S$. In this case, time evolution of the state of $S$ in step (V) is given by (for its derivation, see Appendix B)
@@
for the variable $W$ of an amount of single-shot workhhhSingle-shot work is deterministic and has no fluctuations. on $S$ [23, 27] and the controlled redefined Hamiltonian ^$\widehat{{\cal H}}^{S}_{H,\lambda_{t}}$ of $S$ in the Heisenberg picture with a control parameter $\lambda_{t}$, which is further parameterized by time $t$ in $[0,t_{f}]$ according to the quantum force protocol [23]. Here, ${\cal T}_{>}$ is the time ordering of products and we set $\beta=1/(k_{B}T)$, $Z_{\lambda}={\rm tr}_{S}\bigl{[}e^{-\beta\widehat{{\cal H}}_{H,\lambda}^{S}}\bigr{]}$, and $t_{n}=nt_{f}/N$. This equality is
@@
    The operator to be averaged in the quantum Jarzynski equality (21) is defined over $t_{f}\geq t\geq t^{(0)}$ at an arbitrary intermediate time $t^{(0)}$ in the whole process $[0,t_{f}]$.iiiFor its derivation, see Appendix C.1 of Ref. [10]. Thus, the operator to be averaged is redefined by multiplying it by the factor $e^{-3\sigma}$.
@@
In quantum thermodynamics, the quantum work done by a work reservoir to a thermodynamic quantum mechanical systemjjjIn modern thermodynamics, the thermodynamic systems and the systems in the thermodynamic limit are distinct classes of systems: a microscopic system in thermal contact with a large heat bath is a thermodynamic system. and the heat absorption from a heat bath by the system are distinguished as the change of the energy eigenvalues (the quantum Hamiltonian) of the system and the change of the corresponding occupation probabilities of the system, respectively, in the internal energy of the system [26, 29, 30].
@@ 5 Conclusion
Finally, to conclude this paper, we explain how this result can be positioned within the existing literature on quantum thermodynamics from the perspective of the information-theoretical origins of work, in the setting of the quantum Jarzynski equality, in terms of the density matrix. The explanation consists of three stages.
@@
for the average quantum work $\overline{W}$ done by a work reservoir to a thermodynamic quantum mechanical system, which is in thermal contact with a large heat bath at temperature $T$, and the change in the non-equilibrium free energy $\Delta F$ of the systemkkkNon-equilibrium free energy is obtained from the equilibrium free energy by replacing the thermodynamic entropy with information-theoretical entropy [31]. induced by the work in $\overline{W}$. In general, the work $W$ fluctuates thermally and quantum mechanically. In the thermodynamic limit of the system, the thermal fluctuations in the system become negligible. When the change of the Hamiltonian of the system is cyclic, the change in the non-equilibrium free energy $\Delta F$ of the system has an information-theoretical origin in terms of the density matrix:
@@
where $\widehat{\varrho}$ is the terminal quantum state of the system and $\widehat{\varrho}_{\rm can}$ is the initially prepared canonical thermal equilibrium quantum state of the system at temperature $T$ [32]. Here, the relative von Neumann entropy is always non-negative: $\Delta F\geq 0$. It should be noted that this information-theoretical expression (29) and (30) for the work was extended, in the single-shot work, to the Gibbs-preserving and completely-positive and trace-preserving map of the system in the context of quantum resource theory, where the extractable work from a resource and the work required to create a resource are each individually regarded as the work and the thermodynamic limit of the system is unnecessary [31, 33, 34, 35, 36].
@@
Here, $W_{\rm e.r.}$ is internal work in the composite system of systems $S$ and $M$. Namely, in the composite system, the work for $S$ and that for $M$ cancel each other out and there is no net extractable work and no net energy cost.
We now conclude this paper. As seen from the above explanations of the information-theoretical origins of work in quantum thermodynamics in three stages, the extractable quantum work $-W_{\rm f.b.}$ characterizes the information-to-energy conversion (see Ref. [40] for a renowned experiment in the classical mechanical context), but the classical single-shot work $W_{\rm e.r.}$ characterizes the physicality of the projection hypothesis in the setting of the quantum Jarzynski equality. Namely, in the ensemble interpretation of quantum mechanics, we have theoretically shown a way to experimentally examine the physicality of the projection hypothesis [13]. This is the implication of the main results in the fields of quantum information and quantum thermodynamics.
@@ Appendix C Experimental Platform
In this appendix, we clarify the experimental platform for realizing the four systems which appear in the quantum thermodynamic scheme proposed in Sec. 4. A model of the quantum measurement steps in the proposed scheme is shown in Fig. C.1.
Figure C.1: Quantum measurement steps in the proposed quantum thermodynamic scheme. There are three steps: first, the non-selective measurement (n.s.m.) in the composite system $S=S_{0}+A^{\prime}$; second, the entangling (ent.) process between systems $S_{0}$ in $S$ and $\psi$ in $M$; and third, the von Neumann-type interaction (v.N. int.) between systems $\psi$ and $A$. Each dashed arrow from system $X$ to system $Y$ indicates the interaction process (e.g., the von Neumann-type interaction) between systems $X$ and $Y$, where the state of system $X$ does not change.
In Fig. C.1, $S_{0}$ and $\psi$ are quantum mechanical measured and event-reading systems, respectively. In the context of quantum measurement theory [9], they are distinguished from each other. See also remark (E).